# Graph Transformers: A Survey

### v1

@@ Abstract
Index Terms:
Graph transformer, attention, graph neural network, representation learning, graph learning, network embedding
@@ I Introduction
Graphs, as data structures with high expressiveness, are widely used to present complex data in diverse domains, such as social media, knowledge graphs, biology, chemistry, and transportation networks [1]. They capture both structural and semantic information from data, facilitating various tasks, such as recommendation [2], question answering [3], anomaly detection [4], sentiment analysis [5], text generation [6], and information retrieval [7]. To effectively deal with graph-structured data, researchers have developed various graph learning models, such as graph neural networks (GNNs), learning meaningful representations of nodes, edges and graphs [8]. Particularly, GNNs following the message-passing framework iteratively aggregate neighboring information and update node representations, leading to impressive performance on various graph-based tasks [9]. Applications ranging from information extraction to recommender systems have benefited from GNN modelling of knowledge graphs [10].
@@
More recently, the graph transformer, as a newly arisen and potent graph learning method, has attracted great attention in both academic and industrial communities [11, 12]. Graph transformer research is inspired by the success of transformers in natural language processing (NLP) [13] and computer vision (CV) [14], coupled with the demonstrated value of GNNs. Graph transformers incorporate graph inductive bias (e.g., prior knowledge or assumptions about graph properties) to effectively process graph data [15]. Furthermore, they can adapt to dynamic and heterogeneous graphs, leveraging both node and edge features and attributes. [16]. Various adaptations and expansions of graph transformers have shown their superiority in tackling diverse challenges of graph learning, such as large-scale graph processing [17]. Furthermore, graph transformers have been successfully employed in various domains and applications, demonstrating their effectiveness and versatility.
@@
Existing surveys do not adequately cover the latest advancements and comprehensive applications of graph transformers. In addition, most do not provide a systematic taxonomy of graph transformer models. For instance, Chen et al.  [18] focused primarily on the utilization of GNNs and graph transformers in CV, but they failed to summarize the taxonomy of graph transformer models and ignored other domains, such as NLP. Similarly, Müller et al. [12] offered an overview of graph transformers and their theoretical properties, but they did not provide a comprehensive review of existing methods or evaluate their performance on various tasks. Lastly, Min et al. [19] concentrated on the architectural design aspect of graph transformers, offering a systematic evaluation of different components on different graph benchmarks, but they did not include significant applications of graph transformers or discuss open issues in this field.
@@
    We present a novel taxonomy of graph transformers based on their depth, scalability, and pre-training strategy. We also provide a guide to choosing effective graph transformer architectures for different tasks and scenarios.
@@
An overview of this paper is depicted in Figure 1. The subsequent survey is structured as follows: Section II introduces notations and preliminaries pertaining to graphs and transformers. Section III delves into the design perspectives of graph transformers that encompass graph inductive bias and graph attention mechanisms. Section IV presents a taxonomy of graph transformers categorizing them based on their depth, scalability and pre-training strategy. Additionally, a guide is provided for selecting appropriate graph transformer models for diverse tasks and domains. Section V explores the application perspectives of graph transformers on various node-level, edge-level, and graph-level tasks, along with other application scenarios. Section VI identifies open issues and future directions for research on graph transformers. Lastly, Section VII concludes the paper and highlights its main contributions.
@@ II Notations and Preliminaries
In this section, we present fundamental notations and concepts utilized throughout this survey paper. Additionally, we provide a concise summary of current methods for graph learning and self-attention mechanisms which serve as the basis for graph transformers. Table I includes mathematical notations used in this paper.
@@
[…]
@@ II-A Graphs and Graph Neural Networks
A graph is a data structure consisting of a set of nodes (or vertices) $V$ and a set of edges (or links) $E$ that connect pairs of nodes. Formally, a graph can be defined as $G=(V,E)$, where $V=\{v_{1},v_{2},\dots,v_{N}\}$ is node set with $N$ nodes and $E=\{e_{1},e_{2},\dots,e_{M}\}$ is edge set with $M$ edges. Edge $e_{k}=(v_{i},v_{j})$ indicates the connection between node $v_{i}$ and node $v_{j}$, where $i,j\in\{1,2,\dots,N\}$ and $k\in\{1,2,\dots,M\}$. A graph can be represented by an adjacency matrix $\textbf{A}\in\mathbb{R}^{N\times N}$, where $A_{ij}$ indicates the presence or absence of an edge between node $v_{i}$ and node $v_{j}$. Alternatively, a graph can be represented by the edge list $E\in\mathbb{R}^{M\times 2}$, where each row of $E$ contains the indices of two nodes connected by an edge. A graph can also have node features and edge features that describe the attributes or properties of nodes and edges, respectively. The features of the nodes can be represented by a feature matrix $\textbf{X}\in\mathbb{R}^{N\times d_{n}}$, where $d_{n}$ is the dimension of the node features. The edge features can be represented by a feature tensor $\textbf{F}\in\mathbb{R}^{M\times d_{e}}$, where $d_{e}$ is the dimension of the edge features [20].
Graph learning refers to the task of acquiring low-dimensional vector representations, also known as embeddings for nodes, edges, or the entire graph. These embeddings are designed to capture both structural and semantic information of the graph. GNNs are a type of neural network model that excels at learning from graph-structured data. They achieve this by propagating information along edges and aggregating information from neighboring nodes [21]. GNNs can be categorized into two main groups: spectral methods and spatial methods.
Spectral methods are based on graph signal processing and graph Fourier transform, implementing convolution operations on graphs in the spectral domain [22]. The Fourier graph transform is defined as $\mathbf{\hat{X}}=\textbf{U}^{T}\textbf{X}\textbf{U}$, where $\mathbf{\hat{X}}$ is the spectral representation of node feature matrix X and U is the eigenvector matrix of the normalized graph Laplacian matrix $\textbf{L}=\textbf{I}_{N}-\textbf{D}^{-1/2}\textbf{A}\textbf{D}^{-1/2}$, where $\textbf{I}_{N}$ is the identity matrix and D is diagonal degree matrix with $D_{ii}=\sum_{j=1}^{N}A_{ij}$. Spectral methods can capture global information about the graph, but they suffer from high computational complexity, poor scalability, and the lack of generalization to unseen graphs [23].
@@
$\textbf{h}_{v}^{(l+1)}=\phi\left(\textbf{h}_{v}^{(l)},\bigoplus_{u\in\mathcal{N}(v)}f(\textbf{h}_{u}^{(l)},\textbf{h}_{v}^{(l)},\textbf{e}_{uv})\right)\end{aligned},$ (1)
where $\textbf{h}_{v}^{(l)}$ is the hidden state of node $v$ at layer $l$, $\phi$ is an update function and $\oplus$ is an aggregation function. $\mathcal{N}(v)$ is the set of neighbors of node $v$ and $f$ is a message function that depends on node states and edge features. $\textbf{e}_{uv}$ is the feature vector of the edge between nodes $u$ and $v$. Spatial methods can capture local information of the graph, but they have limitations in modeling long-range dependencies, complex interactions and heterogeneous structures [25].
@@ II-B Self-attention and transformers
Self-attention is a mechanism that enables a model to learn to focus on pertinent sections of input or output sequences [26]. It calculates a weighted sum of all elements in a sequence with weights determined by the similarity between each element and a query vector. Formally, self-attention is defined as:
$\text{Attention}(\textbf{Q},\textbf{K},\textbf{V})=\text{softmax}\left(\frac{\textbf{QK}^{T}}{\sqrt{d_{k}}}\right)\textbf{V}\end{aligned},$ (2)
where Q, K and V are query, key and value matrices,respectively. $d_{k}$ is the dimension of the query and key matrices. Self-attention can capture long-range dependencies, global context and variable-length sequences without using recurrence or convolution.
@@
Transformers are neural network models that use self-attention as the main building block [13]. Transformers consist of two main components: an encoder and a decoder. The encoder takes an input sequence $\textbf{X}=\{\textbf{x}_{1},\textbf{x}_{2},\dots,\textbf{x}_{N}\}$ and generates a sequence of hidden states $\textbf{Z}=\{\textbf{z}_{1},\textbf{z}_{2},\dots,\textbf{z}_{N}\}$. The decoder takes an output sequence $\textbf{Y}=\{\textbf{y}_{1},\textbf{y}_{2},\dots,\textbf{y}_{N}\}$ and generates a sequence of hidden states $\textbf{S}=\{\textbf{s}_{1},\textbf{s}_{2},\dots,\textbf{s}_{N}\}$. The decoder also uses an attention mechanism to attend to the encoder's hidden states. Formally, the encoder and decoder are defined as:
$\textbf{z}_{i}$ $=\text{EncoderLayer}(\textbf{x}_{i},\textbf{Z}_{<i}),$ (3)
$\textbf{s}_{j}$ $=\text{DecoderLayer}(\textbf{y}_{j},\textbf{S}_{<j},\textbf{Z}).$
Here, EncoderLayer and DecoderLayer are composed of multiple self-attention and feed-forward sublayers. Transformers can achieve state-of-the-art results on various tasks, such as machine translation [27], text mining [28, 29], document comprehending [30, 31], image retrieval [32], visual question answering [3], image generation [33], and recommendation systems [34, 35]. An overview of vanilla transformer is shown in Figure 2.
Graph transformers integrate graph inductive bias into transformer to acquire knowledge from graph-structured data [36]. By employing self-attention mechanisms on nodes and edges, graph transformers can effectively capture both local and global information of the graph. In particular, graph transformers exhibit the ability to handle heterogeneous graphs containing diverse types of nodes and edges, as well as complex graphs featuring higher-order structures [12, 37].
@@ III Design Perspectives of graph transformers
In this section, we discuss the primary architectures of graph transformers, aiming to explore their design perspectives in depth. Particularly, we will focus on two key components: graph inductive biases and graph attention mechanisms, to understand how these elements shape graph transformer models’ capabilities.
@@ III-A Graph Inductive Bias
Unlike Euclidean data, such as texts and images, graph data is non-Euclidean data, which has intricate structures and lacks a fixed order and dimensionality, posing difficulties in directly applying standard transformers on graph data [16]. To address this issue, graph transformers incorporate graph inductive bias to encode the structural information of graphs and achieve effective generalization of transformers across new tasks and domains. In this section, we explore the design perspectives of graph transformers through the lens of graph inductive bias. We classify graph inductive bias into four categories: node positional bias, edge structural bias, message-passing bias, and attention bias (see Figure. 3).
Figure 3: Graph Inductive bias: (A) and (B) are local and global node positional biases, respectively. (C) and (D) are local and global edge structural biases, respectively. (E) and (F) are local and global attention biases, respectively. Notably, message-passing bias is a kind of local attention bias.
@@ III-A1 Node Positional Bias
Node positional bias is a crucial inductive bias for graph transformers because it provides information about the relative or absolute positions of nodes in a graph [38]. Formally, given a graph $G=(V,E)$ with $N$ nodes and $M$ edges, each node $v_{i}\in V$ has a feature vector $\textbf{x}_{i}\in\mathbb{R}^{d}_{n}$. A graph transformer aims to learn a new feature vector $\textbf{h}_{i}\in\mathbb{R}^{d}_{k}$ for each node by applying a series of self-attention layers. A self-attention layer can be defined as:
$\textbf{h}_{i}=\sum_{j=1}^{n}a_{ij}W\textbf{x}_{j}+b,$ (4)
where $a_{ij}$ is the attention score between nodes $v_{i}$ and $v_{j}$, measureing the relevance or similarity of their features. $W$ and $b$ are learnable parameters. However this self-attention mechanism does not consider the structural and positional information of nodes, which is crucial for capturing graph semantics and inductive biases [39]. Node positional encoding is a way to address this challenge by providing additional positional features to nodes, reflecting their relative or absolute positions in the graph.
@@
Local Node Positional Encodings. Building on the success of relative positional encodings in NLP, graph transformers leverage a similar concept for local node positional encodings. In NLP, each token receives a feature vector that captures its relative position and relationship to other words in the sequence [40]. Likewise, graph transformers assign feature vectors to nodes based on their distance and relationships with other nodes in the graph [41]. This encoding technique aims to preserve the local connectivity and neighborhood information of nodes, which is critical for tasks like node classification, link prediction, and graph generation.
A proficient approach for integrating local node positional information involves the utilization of one-hot vectors. These vectors represent the hop distance between a node and its neighboring nodes [11]:
$\textbf{p}_{i}=[I(d(i,j)=1),I(d(i,j)=2),...,I(d(i,j)=max)].$ (5)
In this equation, $d(i,j)$ represents the shortest path distance between node $v_{i}$ and node $v_{j}$ and $I$ is an indicator function that returns 1 if its argument is true and 0 otherwise. The maximum hop distance is denoted by $max$. This encoding technique was utilized by Velickovic et al. [42] to enhance their Graph Attention Networks (GATs) with relative position-aware self-attention. Another way to incorporate local node positional encodings in a graph is by using learnable embeddings that capture the relationship between two nodes [43, 44]. This approach is particularly useful when a node has multiple neighbors with different edge types or labels. In this case, the local node positional encoding can be learned based on these edge features:
$\textbf{p}_{i}=[f(\textbf{e}_{ij_{1}}),f(\textbf{e}_{ij_{2}}),...,f(\textbf{e}_{ij_{l}})],$ (6)
where $\textbf{e}_{ij}$ is edge feature between node $v_{i}$ and node $v_{j}$, $f$ is a learnable function that maps edge features to embeddings and $l$ is the number of neighbors considered.
To enhance the implementation of local node positional encodings in a broader context, a viable strategy is leveraging graph kernels or similarity functions to evaluate the structural similarity between two nodes within the graph [45, 37]. For instance, when a node exhibits three neighbors with unique subgraph patterns or motifs in their vicinity, its local node positional encoding can be computed as a vector of kernel values between the node and its neighboring nodes:
$\textbf{p}_{i}=[K(G_{i},G_{j_{1}}),K(G_{i},G_{j_{2}}),...,K(G_{i},G_{j_{l}})].$ (7)
In this equation, $G_{i}$ refers to the subgraph formed by node $v_{i}$ and its neighbors. The function $K$ is a graph kernel function that measures the similarity between two subgraphs. Mialon et al. [46] utilized this approach in their GraphiT model, which incorporates positive definite kernels on graphs as relative positional encodings for graph transformers.
Local node positional encodings enhance the self-attention mechanism of graph transformers by integrating structural and topological information of nodes. This encoding approach offers the advantage of preserving the sparsity and locality of graph structure, resulting in enhanced efficiency and interpretability. However, a limitation of this method is its restricted ability to capture long-range dependencies or global properties of graphs, which are essential for tasks such as graph matching or alignment.
Global Node Positional Encodings. The concept of global node positional encodings is inspired by the use of absolute positional encodings in NLP [47]. In NLP, as mentioned previously, every token receives a feature vector indicating its position within a sequence. Extending this idea to graph transformers, each node can be assigned a feature vector representing its position within the embedding space of the graph [48]. The objective of this encoding technique is to encapsulate the overall geometry and spectrum of graphs, thereby unveiling its intrinsic properties and characteristics.
One method for obtaining global node positional encodings is to leverage eigenvectors or eigenvalues of a matrix representation, such as adjacency matrix or Laplacian matrix [39]. For instance, if a node’s coordinates lie within the first k eigenvectors of graph Laplacian, its global node positional encoding can be represented by the coordinate vector:
$\textbf{p}_{i}=[u_{i1},u_{i2},...,u_{ik}],$ (8)
where $u_{ij}$ is $j$-th component of $i$-th eigenvector of the graph Laplacian matrix. One alternative approach to incorporating global node positional encodings is by utilizing diffusion or random walk techniques, such as personalized PageRank or heat kernel [49]. For example, if a node possesses a probability distribution over all other nodes in the graph, following a random walk [50], its global node positional encoding can be represented by this probability vector:
$\textbf{p}_{i}=[\pi_{i1},\pi_{i2},...,\pi_{iN}],$ (9)
where $\pi_{ij}$ is the probability of reaching node $v_{j}$ from node $v_{i}$ after performing a random walk on the graph.
A more prevalent approach to implementing global node positional encodings is to utilize graph embedding or dimensionality reduction techniques that map nodes to a lower-dimensional space while maintaining a sense of similarity or distance [51]. For instance, if a node possesses coordinates in a space derived from applying multi-dimensional scaling or graph neural networks to the graph, its global node positional encoding can be represented by that coordinate vector:
$\textbf{p}_{i}=[y_{i1},y_{i2},...,y_{ik}],$ (10)
where $y_{ij}$ is $j$-th component of $i$-th node embedding in a $k$-dimensional space, which can be obtained by minimizing the objective function that preserves graph structure:
$\min_{\textbf{Y}}\sum_{i,j=1}^{N}\textbf{w}_{ij}\|\textbf{y}_{i}-\textbf{y}_{j}\|^{2}.$ (11)
Here, $\textbf{w}_{ij}$ is a weight matrix that reflects the similarity or distance between node $v_{i}$ and node $v_{j}$, $Y$ is the matrix of node embeddings.
The primary aim of global node positional encodings is to improve the representation of node attributes in graph transformers by incorporating geometric and spectral information from graphs. This encoding method offers the advantage of capturing long-range dependencies and overall graph characteristics, benefiting tasks like graph matching and alignment. However, a drawback of this encoding approach is that it could undermine the sparsity and locality of graph structures, potentially impacting efficiency and interpretability.

### v2

@@ Abstract
publicationid: pubid: 0000–0000/00\$00.00 © 2025 IEEE
@@ I Introduction
Graphs, as data structures with high expressiveness, are widely used to present complex data in various domains, such as social networks, knowledge graphs, biology, chemistry, and transportation networks [1]. They capture both structural and semantic information from data, facilitating various tasks, such as recommendation [2], question answering [3], anomaly detection [4], sentiment analysis [5], text generation [6], and information retrieval [7]. To effectively deal with graph-structured data, researchers have developed various graph learning models, such as graph neural networks (GNNs), learning meaningful representations of nodes, edges and graphs [8]. In particular, GNNs following the message-passing framework iteratively aggregate neighboring information and update node representations, leading to impressive performance on various graph-based tasks [9]. Applications ranging from information extraction to recommender systems have benefited from GNN modelling of knowledge graphs [10].
@@
More recently, the graph transformer, as a new and powerful graph learning method, has attracted great attention in both the academic and industrial communities [11, 12]. Graph transformer research is inspired by the success of transformers in natural language processing (NLP) [13] and computer vision (CV) [14], coupled with the demonstrated value of GNNs. While GNNs rely on local message passing, which limits their ability to capture long-range dependencies and global structure, graph transformers overcome this limitation by using self-attention mechanisms that model interactions across the entire graph. Graph transformers incorporate graph inductive bias (e.g., prior knowledge or assumptions about graph properties) to effectively process graph data [15]. Furthermore, they can adapt to dynamic and heterogeneous graphs, leveraging both node and edge features and attributes [16]. However, this expressiveness comes at the cost of higher computational complexity and challenges in scaling to large graphs. Thus, a key trade-off in using graph transformers lies between modeling power and computational efficiency. Various adaptations and expansions of graph transformers have shown their superiority in tackling diverse challenges of graph learning, such as large-scale graph processing [17]. Furthermore, graph transformers have been successfully employed in various domains and applications, demonstrating their effectiveness and versatility.
@@
Existing surveys do not adequately cover the latest advancements and comprehensive applications of graph transformers. In addition, most do not provide a systematic taxonomy of graph transformer models. For instance, Chen et al. [18] focused primarily on the utilization of GNNs and graph transformers in CV, but they failed to summarize the taxonomy of graph transformer models and ignored other domains, such as NLP. Similarly, Müller et al. [12] offered an overview of graph transformers and their theoretical properties, but they did not provide a comprehensive review of existing methods or evaluate their performance on various tasks. Lastly, Min et al. [19] concentrated on the architectural design aspects of graph transformers, offering a systematic evaluation of different components on different graph benchmarks, but did not include significant applications of graph transformers or discuss open issues in this field.
@@
    We present a novel taxonomy of graph transformers based on their depth, scalability, and pre-training strategy. We also provide a guide for choosing effective graph transformer architectures for different tasks and scenarios.
@@
An overview of this paper is depicted in Figure 1. The subsequent survey is structured as follows: Section II introduces notation and preliminaries pertaining to graphs and transformers. Section III delves into the design perspectives of graph transformers that encompass graph inductive bias and graph attention mechanisms. Section IV presents a taxonomy of graph transformers categorizing them based on their depth, scalability and pre-training strategy. In addition, a guide is provided to select appropriate graph transformer models for various tasks and domains. Section V explores the application perspectives of graph transformers in various node-level, edge-level, and graph-level tasks, along with other application scenarios. Section VI provides a guide for designing effective graph transformer architectures for different tasks and scenarios. Section VII identifies open issues and future directions for research on graph transformers. Lastly, Section VIII concludes the paper and highlights its main contributions.
@@ II Notations and Preliminaries
In this section, we present the fundamental notations and concepts used throughout the survey. Additionally, we summarize current methods for graph learning and self-attention mechanisms, which form the foundation for graph transformers. The notations listed in Table I provide the mathematical language used in subsequent sections, including definitions of graphs, node and edge features, attention parameters, and positional encodings. These elements are essential to describe core mechanisms of graph neural networks (Section III), attention computation in graph transformers (Section IV), and task-specific architectures (Section V).
@@
$\textbf{Q},\textbf{K},\textbf{V}$ Query, key and value matrices for self-attention
@@ II-A Graphs and Graph Neural Networks
A graph is a data structure consisting of a set of nodes (or vertices) $V$ and a set of edges (or links) $E$ connecting pairs of nodes. Formally, a graph is defined as $G=(V,E)$, where $V=\{v_{1},v_{2},\dots,v_{N}\}$ is the set of nodes with $N$ nodes and $E=\{e_{1},e_{2},\dots,e_{M}\}$ is the set of edges with $M$ edges. The edge $e_{k}=(v_{i},v_{j})$ connects node $v_{i}$ and node $v_{j}$, with $i,j\in\{1,2,\dots,N\}$ and $k\in\{1,2,\dots,M\}$. A graph can be represented by an adjacency matrix $\textbf{A}\in\mathbb{R}^{N\times N}$, where $A_{ij}$ indicates the presence or absence of an edge between nodes $v_{i}$ and $v_{j}$. Alternatively, it can be represented by the edge list $E\in\mathbb{R}^{M\times 2}$, where each row of $E$ contains the indices of two connected nodes. Graphs can also include node features and edge features, represented by a node feature matrix $\textbf{X}\in\mathbb{R}^{N\times d_{n}}$ (where $d_{n}$ is the dimension of node features) and an edge feature tensor $\textbf{F}\in\mathbb{R}^{M\times d_{e}}$ (where $d_{e}$ is the dimension of edge features) [20].
Graph learning involves deriving low-dimensional vector representations, or embeddings, for nodes, edges, or entire graphs to capture structural and semantic information. For instance, embeddings represent users in social networks by encoding interactions; capture atom types and chemical bonds in molecular graphs for drug discovery; and characterize intersections and roads in navigation systems for efficient route planning. Graph neural networks (GNNs) effectively learn from graph-structured data by propagating and aggregating information among neighboring nodes [21]. GNNs are primarily categorized into spectral and spatial methods. Spectral methods utilize graph signal processing to perform convolution operations in the spectral domain using the Fourier graph transform defined as $\mathbf{\hat{X}}=\textbf{U}^{T}\textbf{X}\textbf{U}$, where $\mathbf{U}$ contains eigenvectors of the normalized graph Laplacian $\mathbf{L}=\mathbf{I}_{N}-\mathbf{D}^{-1/2}\mathbf{A}\mathbf{D}^{-1/2}$ [22]. While spectral methods capture global graph properties, they exhibit high computational complexity and limited scalability [23].
@@
$\begin{aligned} \textbf{h}_{v}^{(l+1)}=\phi\left(\textbf{h}_{v}^{(l)},\bigoplus_{u\in\mathcal{N}(v)}f(\textbf{h}_{u}^{(l)},\textbf{h}_{v}^{(l)},\textbf{e}_{uv})\right)\end{aligned},$ (1)
where $\textbf{h}_{v}^{(l)}$ is the hidden state of node $v$ at layer $l$, $\phi$ is an update function and $\oplus$ is an aggregation function. $\mathcal{N}(v)$ is the set of neighbors of node $v$ and $f$ is a message function that depends on the states of the nodes and the edge features. $\textbf{e}_{uv}$ is the edge feature vector between nodes $u$ and $v$. Spatial methods can capture local information on the graph, but have limitations in modeling long-range dependencies, complex interactions, and heterogeneous structures [24].
@@ II-B Self-attention and Transformers
Self-attention is a core mechanism that enables models to relate different parts of the input sequence to each other [25]. Intuitively, it helps the model decide how much attention to pay to other elements when processing a particular item. For example, in a sentence, understanding the meaning of a word often depends on its surrounding words; self-attention allows the model to capture such contextual relationships. Formally, self-attention computes a weighted sum of all elements in a sequence, where the weights reflect the similarity between each element and a query vector. It is defined as:
$\begin{aligned} \text{Attention}(\textbf{Q},\textbf{K},\textbf{V})=\text{softmax}\left(\frac{\textbf{QK}^{T}}{\sqrt{d_{k}}}\right)\textbf{V}\end{aligned},$ (2)
where Q, K and V are query, key and value matrices, respectively. $d_{k}$ is the dimension of the query and key matrices. Self-attention can capture long-range dependencies, global context, and variable-length sequences without using recurrence or convolution.
Transformers are neural network models that use self-attention as the main building block [13]. Transformers consist of two main components: an encoder and a decoder. The encoder takes an input sequence $\textbf{X}=\{\textbf{x}_{1},\textbf{x}_{2},\dots,\textbf{x}_{N}\}$ and generates a sequence of hidden states $\textbf{Z}=\{\textbf{z}_{1},\textbf{z}_{2},\dots,\textbf{z}_{N}\}$. The decoder takes an output sequence $\textbf{Y}=\{\textbf{y}_{1},\textbf{y}_{2},\dots,\textbf{y}_{N}\}$ and generates a sequence of hidden states $\textbf{S}=\{\textbf{s}_{1},\textbf{s}_{2},\dots,\textbf{s}_{N}\}$. In the final step of the decoder, these hidden states are passed through a softmax layer to compute probabilities over possible output tokens, which can then be used for downstream tasks such as sequence generation or classification. The decoder also uses an attention mechanism to attend to the encoder’s hidden states. Formally, the encoder and decoder are defined as: $\textbf{z}_{i}=\text{EncoderLayer}(\textbf{x}_{i},\textbf{Z}_{<i})$, $\textbf{s}_{j}=\text{DecoderLayer}(\textbf{y}_{j},\textbf{S}_{<j},\textbf{Z}).$ Here, EncoderLayer and DecoderLayer are composed of multiple self-attention and feed-forward sublayers. Transformers can achieve state-of-the-art results in various tasks, such as machine translation [26], text mining [27], document comprehension [28], image retrieval [29], visual question answering [3], and image generation [30]. An overview of the vanilla transformer is shown in Figure 2.
Graph transformers integrate graph inductive bias into transformers to acquire knowledge from graph-structured data [31]. By employing self-attention mechanisms on nodes and edges, graph transformers can effectively capture both local and global information of the graph. However, traditional transformer models, which are designed for sequence data, cannot be directly applied to graph data because graph-structured data does not have a natural ordering, as is the case with sequence data. In graph data, nodes and edges are interconnected in a non-sequential manner, and this necessitates the modification of the standard self-attention mechanism. To address this challenge, graph transformers modify the attention mechanism by incorporating graph-specific inductive biases such as positional encoding based on graph structure and attention over both node and edge features. These adaptations enable graph transformers to learn dependencies not only along the sequential dimensions but also in the graph topology, capturing relationships between nodes even in non-sequential configurations. [12, 32].
@@ III Design Perspectives of Graph Transformers
In this section, we introduce the concept of “design perspectives” in Graph Transformers, referring to the core principles and strategies that guide how graph-specific inductive biases, attention mechanisms, and structural properties are integrated into the Transformer architecture. By examining these underlying design considerations, we aim to illustrate how Graph Transformers can effectively capture both local and global relationships in graph-structured data. While this section focuses on fundamental design concepts, Section IV provides a detailed taxonomy of Graph Transformer architectures, covering the diverse ways in which these design principles are operationalized to address various tasks and trade-offs in graph learning.
This section examines how graph inductive biases and graph attention mechanisms are incorporated into graph transformer models to handle structured data effectively.
@@ III-A Graph Inductive Bias
Unlike Euclidean data such as text and images, graph data is non-Euclidean with complex structures and no fixed order or dimensionality, making it challenging to apply standard transformers [16]. Graph transformers address this issue by integrating graph inductive bias, which encodes structural information for better generalization. We classify graph inductive bias in graph transformers into four categories: node positional bias, edge structural bias, message-passing bias, and attention bias.
@@ III-A1 Node Positional Bias
Node positional bias provides relative or absolute positional information for nodes in a graph [33]. Given a graph $G=(V,E)$ with $N$ nodes and $M$ edges, each node $v_{i}\in V$ has a feature vector $\textbf{x}_{i}\in\mathbb{R}^{d}_{n}$. A graph transformer aims to learn a new embedding vector $\textbf{h}_{i}\in\mathbb{R}^{d}_{k}$ for each node by leveraging self-attention layers during training. A self-attention layer is defined as: $\textbf{h}_{i}=\sum_{j=1}^{n}a_{ij}W\textbf{x}_{j}+b,$ where $a_{ij}$ is the attention score between nodes $v_{i}$ and $v_{j}$, measuring the relevance or similarity of their features, and $W$ and $b$ are learnable parameters. Because this mechanism initially lacks structural and positional information from the nodes, node positional encoding addresses these gaps by adding positional features that capture graph semantics and inductive biases [34].
@@
Local Node Positional encoding. Building on the success of relative positional encoding in NLP, graph transformers adopt a similar concept for local node positional encoding. In NLP, each token receives a feature vector capturing its relative position and relationship to other words [35]. Similarly, graph transformers assign feature vectors to nodes based on their distance and relationships with other nodes [36], preserving local connectivity and neighborhood information, which benefits tasks such as node classification, link prediction, and graph generation. A frequently used approach leverages one-hot vectors representing the hop distance between a node and its neighbors [11]: $\textbf{p}_{i}=[I(d(i,j)=1),I(d(i,j)=2),...,I(d(i,j)=max)],$ where $d(i,j)$ is the shortest path distance between nodes $v_{i}$ and $v_{j}$, and $I$ is an indicator function. The maximum hop distance is maxmax. This technique was used by Velickovic et al. [37] to enhance Graph Attention Networks (GATs) with position-aware self-attention. Another method involves learnable embeddings capturing the relationship between nodes [38, 39], useful when a node has multiple neighbors with different edge types or labels: $\textbf{p}_{i}=[f(\textbf{e}_{ij_{1}}),f(\textbf{e}_{ij_{2}}),...,f(\textbf{e}_{ij_{l}})],$ where $\textbf{e}_{ij}$ is the edge feature between nodes $v_{i}$ and $v_{j}$, $f$ is a learnable function mapping edge features to embeddings, and $l$ is the number of neighbors considered.
Another option is to leverage graph kernels or similarity functions to measure structural similarity between nodes [40, 32]: $\textbf{p}_{i}=[K(G_{i},G_{j_{1}}),K(G_{i},G_{j_{2}}),...,K(G_{i},G_{j_{l}})],$ where $G_{i}$ is the subgraph consisting of node $v_{i}$ and its neighbors, and $K$ is a graph kernel function measuring subgraph similarity. Mialon et al. [41] used this idea in their GraphiT model, applying positive definite kernels as relative positional encoding. Although local node positional encoding preserves the sparsity and locality of graph structure and can improve efficiency and interpretability, it has limited ability to capture long-range dependencies or global properties of graphs, which are often required for tasks like graph matching or alignment.
Global Node Positional encoding. Global node positional encoding takes inspiration from absolute positional encoding in NLP [42], assigning to each node a characteristic vector that represents its position within the embedding space of the graph [43]. One method for obtaining global node positional encoding leverages eigenvectors or eigenvalues of a matrix representation (e.g., adjacency or Laplacian) [34]: $\textbf{p}_{i}=[u_{i1},u_{i2},...,u_{ik}],$ where $u_{ij}$ is the $j$-th component of the $i$-th eigenvector of the graph Laplacian. An alternative approach relies on diffusion or random walk techniques, such as personalized PageRank or heat kernel [44]. In this case, the global positional encoding of a node can be represented by its random walk transition probabilities [45]: $\textbf{p}_{i}=[\pi_{i1},\pi_{i2},...,\pi_{iN}],$ where $\pi_{ij}$ is the probability of reaching node $v_{j}$ from node $v_{i}$. Another common method is graph embedding or dimensionality reduction, mapping nodes into a lower-dimensional space while preserving similarity or distance [46]. In such cases: $\textbf{p}_{i}=[y_{i1},y_{i2},...,y_{ik}],$ where $y_{ij}$ is the $j$-th component of the $i$-th node embedding, obtained by minimizing an objective function that preserves the graph’s structure: $\min_{\textbf{Y}}\sum_{i,j=1}^{N}\textbf{w}_{ij}\|\textbf{y}_{i}-\textbf{y}_{j}\|^{2}.$ Here, $\textbf{w}_{ij}$ is a weight matrix reflecting node similarity or distance. The primary aim of global node positional encoding is to enhance node attribute representation by incorporating geometric and spectral graph information. This method captures long-range dependencies and overall graph characteristics, aiding tasks like graph matching and alignment. However, it can reduce sparsity and locality in graph representations, potentially affecting efficiency and interpretability.